Match the graph with one of the equations.
Crosses at (0,2) & (2,0)

A) y = –x – 2
B) y = x – 2
C) y = –x + 2
D) y = x + 2

Answer by checkley75(3666)

FIRST WE FIND THE SLOPE
SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(0-2)/(2-0)=-2/2=-1
NOW WE SUBSTITUTE ONE SET OF (X,Y) POINTS & SOLVE FOR b THE Y INTERCEPT.
2=-1*0+b
b=2
THUS THE EQUATION IS [Y=-X+2]
THE ANSWER IS C)